# Investigating the correlation between ZTF TDEs and IceCube high-energy neutrinos

### v2

@@ 5.1 The systematic errors of alert events
WEIGHT+Syst. | 0.48 | 0.51 | 0.49 | All TDEs |
WEIGHT+Syst. | 0.045 | 0.060 | 0.063 | TDEs with IR echoes |
WEIGHT+Delay | 0.23 | 0.23 | 0.20 | All TDEs |
WEIGHT+Delay | 0.0080 | 0.035 | 0.022 | TDEs with IR echoes |
@@ 5.2 Time delay factor
In this analysis, the temporal coincidence criterion is required that the discovery dates of TDEs are earlier than the arrival dates of alert events. In this scenario, we estimate the values of correlation between TDEs/subsample of ZTF-BTS catalog and IceCube neutrinos, and the most significant correlation is given by the TDEs with IR echoes subsample.
@@
After adding the time delay factor, the TS values are mainly contributed by the AT2019dsg event. By realizing 104 times MC simulations, the values for the three MC simulation methods are summarized in Table 5. Using the WEIGHT method, compared to the results without the time delay factor, the best value of the all-TDEs sample decreases to 0.20, while that for the TDEs with IR echoes subsample decreases to 0.0080 (). After introducing the time delay factor, the subsample with IR echoes also presents the most significant correlation with IceCube neutrinos compared to the all-TDEs sample.
@@ 5.3 On the connection between high-energy neutrinos and IR emissions
TDEs have increasingly been discovered through optical transient surveys and have been proposed as potential sources of high-energy neutrinos over the past few years. Interestingly, it is noted that the first three TDE candidates reported to be possibly associated with high-energy neutrinos, including the confirmed TDE AT2019dsg and the candidates AT2019fdr and AT2019aalc (Stein et al., 2021; Reusch et al., 2022; van Velzen et al., 2024), as well as some other similar events reported subsequently (Jiang et al., 2023; Yuan et al., 2024; Li et al., 2024), all exhibited infrared echo emissions. In this work, our analysis also shows that the TDEs with IR flares display a stronger correlation with high-energy neutrinos, which further supports previous findings that the TDEs coinciding with neutrinos tend to have prominent IR echo emissions.
The IR emission typically originates from a dust torus surrounding the supermassive black hole: the the optical-ultraviolet or X-ray radiation from the TDE outburst is absorbed by the dust and re-emitted in the IR band, indicating the presence of a dense environment around such TDEs (Lu et al., 2016; van Velzen et al., 2016; Jiang et al., 2016). Although not fully clear, possible reasons for TDEs with IR echoes are more likely to be associated with IceCube high-energy neutrinos include (Wu et al., 2022; Mou et al., 2021; Winter & Lunardini, 2023; Yuan & Winter, 2023; Yuan et al., 2024) (1) Significant IR echoes imply a dusty/cloudy environment around the black hole, where winds or outflows interacting with the circumnuclear medium can drive the formation of shocks, which accelerate protons to high energies and lead to neutrino production. (2) The dense gas and dust in the dust torus provide target material for the accelerated high-energy protons. Compared to TDEs without a dust torus, such an environment significantly enhances the rate of neutrino production via interactions. (3) The IR photons produced by dust echoes serve as target photons for interactions, and accelerated protons interact with the IR photons, though this requires the protons to be accelerated to sufficiently high energies.
@@ 6 Summary
Currently, the main contributors to the IceCube neutrino flux are still obscure. Among various astrophysical sources, TDEs have been suggested as possible emitters of IceCube neutrinos. Therefore, investigating the correlation between the TDE population and IceCube neutrinos could help us better understand whether TDEs could really serve as neutrino emitters. In this work, we study the correlation between the TDEs/TDE candidates from the ZTF-BTS catalog and the population of IceCube neutrinos. Considering that good spatial localization is crucial for establishing reliable association, we employ two methods to handle the alert events with large error circles. One method involves directly excluding alert events with error area 30 deg2, while another method introduces a weighting factor. We quantify the significance of the correlation using MC simulations. We find that the subsample of TDEs with IR echoes presents the most significant correlation with IceCube neutrinos, though still below the confidence level. We also discuss the case that includes a systematic error in the events’ error radii. Given that only the statistical error is publicly available for alert events, we add a systematic error of , which is the upper limit of the systematic error estimated by Aartsen et al. (2013c), in quadrature to the statistical error. There are no additional associated events arising from the inclusion of systematic error and the obtained results could be treated as the upper limits of the values. Furthermore, for the potentially associated events reported so far, the arrival dates of neutrinos are within an interval of year after the discovery dates of the TDEs/TDE candidates. Therefore, we tentatively introduce a time delay () term into the weighting factor in our MC simulation to reduce the weight of large time-delayed events. By combining the time delay factor, the best values decrease to 0.20 and 0.0080 () for the all-TDE sample and the subsample with IR echoes, respectively. However, the lowest value remains below the 3 confidence level, even though trial factors have not been taken into account yet.
@@
In total, at present no significant correlation is found in the population analysis between the ZTF-BTS TDEs and IceCube alert neutrino events. Our work is the first research to quantify the -value of the TDE-neutrino association with actual TDE and neutrino observations. We find that the TDEs with IR flares have a stronger correlation with high-energy neutrinos, consistent with previous findings.
@@
In our sample, only eight TDEs have IR flares and two of them satisfy the spatial-temporal association with high-energy neutrinos. The number of TDEs with IR flares is still limited, which hampers us from obtaining conclusive results. In the future, the persistent optical survey conducted by ZTF-BTS (Perley et al., 2020) is expected to detect more TDEs, in the order of 10 TDEs per year. Moreover, hundreds to thousands of TDEs are expected to be discovered per year by advanced surveys such as the Legacy Survey of Space and Time (LSST) at the Vera Rubin Observatory (Ivezić et al., 2019; Bricman & Gomboc, 2020) and the Wide Field Survey Telescope (WFST, Lin et al. (2022); Wang et al. (2023)). It is also expected that the planned NEO Surveyor (Mainzer et al., 2023), the successor to the NEOWISE mission, will detect more IR flares and help build a larger sample of TDEs with IR echoes, addressing the current issue of too small sample size. By identifying more cases of associations between IR-TDEs and IceCube high-energy neutrinos, we will hopefully further uncover the intrinsic connection between the TDE population and high-energy neutrinos, and understand the role that IR echoes play in the neutrino emission process.

### v3

@@ 4.2 The systematic errors of alert events
WEIGHT+Syst. | 0.48 | 0.50 | 0.49 | All TDEs |
WEIGHT+Syst. | 0.046 | 0.057 | 0.056 | TDEs with IR echoes |
WEIGHT+Delay | 0.22 | 0.22 | 0.20 | All TDEs |
WEIGHT+Delay | 0.0077 | 0.027 | 0.020 | TDEs with IR echoes |
@@ 4.3 Time delay factor
In the above analysis, the temporal coincidence criterion only requires that the discovery dates of TDEs are earlier than the arrival dates of alert events. Under this criterion, we estimate the values of correlation between TDEs/subsample of ZTF-BTS catalog and IceCube neutrinos, and the most significant correlation is given by the subsample of TDEs with IR echoes.
@@
After adding the time delay factor, the TS values are mainly contributed by the AT2019dsg event. By realizing 104 times MC simulations, the values for the three MC simulation methods are summarized in Table 5. Using the WEIGHT method, compared to the results without the time delay factor, the best value of the all-TDEs sample decreases to 0.20, while that for the TDEs with IR echoes subsample decreases to 0.0077 (). After introducing the time delay factor, the subsample with IR echoes also presents the most significant correlation with IceCube neutrinos compared to the all-TDEs sample.
5 Discussion
In Table 6, we summarize the properties of TDEs associated with alert events found in previous works and in this work. We find that most of the listed properties are consistent for all the TDEs. For these TDEs, all exhibit bright IR echoes, and except for the two TDE candidates identified by Jiang et al. (2023) in the IR data, all have been reported in both optical and X-ray observations. Except for AT2019azh, the time differences between IR peaks and arrival dates of alert events range from 18 to 159 days in the source frame. The arrival dates of alert events are closer to the IR peaks than to the optical peaks (148 to 228 days in the source frame except AT2019azh), supporting the connection between IR photons and neutrinos.
@@ 5.1 On the connection between high-energy neutrinos and IR emissions
TDEs have increasingly been discovered through optical transient surveys and have been proposed as potential sources of high-energy neutrinos over the past few years. Interestingly, it is noted that the first three TDE candidates reported to be possibly associated with high-energy neutrinos, including the confirmed TDE AT2019dsg and the candidates AT2019fdr and AT2019aalc (Stein et al., 2021; Reusch et al., 2022; van Velzen et al., 2024), as well as some other similar events reported subsequently (Jiang et al., 2023; Yuan et al., 2024; Li et al., 2024), all have IR echo emission being detected. Here we discuss the possible connection between high-energy neutrinos and IR echo emissions. In this work, our analysis also shows that the TDEs with IR flares display a stronger correlation with high-energy neutrinos, which further supports previous findings that the TDEs coinciding with neutrinos tend to have prominent IR echo emissions.
The IR echo emission typically originates from a dust torus surrounding the supermassive black hole: the optical-ultraviolet or X-ray radiation from the TDE outburst is absorbed by the dust and re-emitted in the IR band, indicating the presence of a dense environment around such TDEs (Lu et al., 2016; van Velzen et al., 2016; Jiang et al., 2016). Although not fully clear, possible reasons for TDEs with IR echoes are more likely to be associated with IceCube high-energy neutrinos have been widely discussed (Wu et al., 2022; Mou et al., 2021; Winter & Lunardini, 2023; Yuan & Winter, 2023; Yuan et al., 2024): (1) Significant IR echoes imply a dusty/cloudy environment around the black hole, where winds or outflows interacting with the circumnuclear medium can drive the formation of shocks, which accelerate protons to high energies and lead to neutrino production. (2) The dense gas and dust in the dust torus themself provide target material for the accelerated high-energy protons. Compared to TDEs without a dust torus, such an environment significantly enhances the rate of neutrino production via interactions. (3) The IR photons produced by dust echoes serve as target photons for interactions, and accelerated protons interact with the IR photons, though this requires the protons to be accelerated to sufficiently high energies.
@@
TDE event | log | log | log | Ref | ||||
(days) | (days) | (erg s-1) | (erg s-1) | (M⊙) | (TeV) | |||
AT2019dsg | 150 | -19 | 43.30 | 43.4 | 0.0512 | 6.74 | 217.42 | (5),(6) |
AT2019fdr | 289 | -75 | 44.56 | 43.1 | 0.267 | 7.10 | 82.186 | (4),(6) |
AT2019aalc | 148 | -91 | 43.80 | 42.1 | 0.0360 | 7.23 | 176.48 | (6) |
AT2021lwx | 370 | 317 | 45.49 | 45.2 | 0.995 | 8.00 | 106.32 | (7),(8) |
SDSSJ1048+1228 | - | -120 | 42.84 | - | 0.0537 | 7.86 | 375.23 | (2) |
SDSSJ1649+2625 | - | -69 | 43.20 | - | 0.0588 | 6.49 | 82.186 | (2) |
AT2019azh | 1556 | 1560 | 42.06 | 43.1 | 0.0223 | 6.50 | 154.28 | This work, (1),(2) |
AT2019azh | 1416 | 1420 | 42.06 | 43.1 | 0.0223 | 6.50 | 54.85 | This work, (1),(2) |
a The arrival date of alert event subtracts the date of TDE optical peak.
b The arrival date of alert event subtracts the date of TDE IR peak.
c The luminosity values are the total of W1 and W2 bands.
d The X-ray luminosity covers an energy range of 0.3-10 keV for AT2021lwx and AT2019azh, while the others have an energy range of 0.2-10 keV.
5.2 The association between AT2019azh and IC230217A
In this work, we find that AT2019azh is spatially associated with two alert events (IC230217A and IC230707B). Among the two, IC230707B has a very large localization uncertainty (10.16 degree radius, corresponding to an area of 323 deg2), which will result in a high probability of coincidental association with TDEs. As shown in Table 3, there are two TDEs (AT2019azh and AT2022bdw) associated with IC230707B within its location error circle, and AT2022bdw is even closer to the nominal position of IC230707B and has a shorter time difference . Considering these, individually discussing such events like IC230707B with poor localization is not particularly meaningful. Therefore, we focus on discussing the association between AT2019azh and the alert IC230217A.
AT2019azh is a thermal TDE exhibiting unusually long-lived radio emission and delayed X-ray brightening (Goodwin et al., 2022). Its optical light curve peaked 27 days after discovery. Radio observations revealed a gradual brightening over years post-disruption, followed by a decline attributed to either interaction with the circumnuclear medium or the shutdown of the central engine (Goodwin et al., 2022). Even 850 days after the TDE, AT2019azh still exhibited strong radio emission. The inferred radio outflow is non-relativistic, plausibly a quasi-spherical or mildly collimated wind from accretion. These radio observations confirm that particle acceleration is occurring. AT2019azh shares many similar characteristics with AT2019dsg (Stein et al., 2021): both are thermal TDEs with sub-relativistic outflows, and their parameters in Table 7 show significant similarity. AT2019azh also displays a rare bright X-ray emission among the TDE population, with a peak luminosity of 1.2 erg/s (Liu et al., 2022; Hinkle et al., 2020), similar to those observed in AT2019dsg, AT2019fdr, and AT2019aalc. The late-time radio behavior of AT2019azh is also similar to that of AT2019dsg (Stein et al., 2021). Both TDEs show initially increasing energy and flux density attributed to constant energy injection from the outflow. These similarities suggest that AT2019azh is also a potential neutrino-emitting source.
However, compared to AT2019dsg, the neutrino event associated with AT2019azh occurred 1500 days after the onset of the TDE, and such a long time delay challenges the possibility that the neutrino originated from AT2019azh. Considering that the IR echo emission from this TDE was only observed at relatively early stages (see Fig. 2), if AT2019azh is indeed associated with IC230217A, then its neutrino production might differ from that of AT2019dsg, AT2019fdr, AT2019aalc (Winter & Lunardini, 2023) and AT2019lwx (Yuan et al., 2024) for which, as discussed in Sec. 5.1, IR photons may serve as the target photons in the neutrino production process.
Although there is a significant delay between IC230217A and the TDE, the radio and X-ray emissions from AT2019azh last for a very long time, making neutrino production still plausible at . The radio signal typically serves as a tracer for the particle acceleration processes required for neutrino generation. The neutrino energy of IC230217A reported by IceCube is 54.9 TeV. This corresponds to parent proton energy of 1 PeV under the typical assumption of a conversion fraction of from parent protons to neutrinos. Based on the velocity and size of the outflow given by the final observational epoch in Table 2 of (Goodwin et al., 2022), the Hillas criterion indicates that the outflow is capable of accelerating protons to . Additionally, to produce neutrinos with an energy of 54.9 TeV, the required target photon energy is and is in the X-ray band. The detection of X-ray emission even at 1500 days post-disruption (Guolo et al., 2024), albeit at low flux, suggests that neutrino production via interactions remains possible. However, the neutrino yield from late-time interactions may be relatively low. Another possible neutrino production mechanism is interactions, where protons in the circumnuclear medium are more likely to serve as target particles in the hadronic processes. For such a channel, interactions between the outflow and surrounding materials (e.g., unbound stellar debris or clouds) form shocks and lead to particle acceleration. The accelerated protons could then interact with also these materials and produce neutrinos (Murase et al., 2020; Wu et al., 2022; Winter & Lunardini, 2023). Neutrinos produced via the process typically have lower energies, which is also consistent with the case of IC230217A.
In all, the optical, radio, and X-ray emissions detected at a late stage suggest activity of the central engine or interactions between the outflow and the circumnuclear medium at around 1500 days, which may generate shocks and accelerate protons, leading to neutrino production. Therefore, despite the large time delay, AT2019azh remains a promising neutrino source candidate.
AT2019azh | AT2019dsg | References | |
[erg/s] | (1), (1) | ||
[erg/s] | (1), (3) | ||
[erg/s] | (2), (2) | ||
[cm] | (2), (2) | ||
(2), (2) | |||
[erg] | (1), (3) | ||
(1), (3) | |||
[G] | (1), (3) | ||
[cm] | (1), (3) | ||
[cm-3] | (1), (1) | ||
Redshift | 0.0223 | 0.0512 | (3), (3) |
a∼c Blackbody luminosity, radius and temperature.
d The energy of the outflow, estimated using radio observations during the period from 132 to 850 days post-disruption for AT2019azh and from 42 to 178 days for AT2019dsg.
f The same as d, but for the outflow velocity in units of .
g For AT2019azh, this row presents the estimated magnetic field during the 132-850 day period post-TDE, while for AT2019dsg it corresponds to the estimation at the arrival date (178 day) of the neutrino.
h The same as g, but for the derived radius of the outflow.
i The same as g, but for the ambient electron density.
@@ 6 Summary
Currently, the main contributors to the IceCube neutrino flux are still obscure. Among various astrophysical sources, TDEs have been suggested as possible emitters of IceCube neutrinos. Therefore, investigating the correlation between the TDE population and IceCube neutrinos could help us better understand whether TDEs could really serve as neutrino emitters. In this work, we study the correlation between the TDEs/TDE candidates from the ZTF-BTS catalog and the population of IceCube neutrinos. Considering that good spatial localization is crucial for establishing a reliable association, we employ two methods to handle the alert events with large error circles. One method involves directly excluding alert events with error area 30 deg2, while another method introduces a weighting factor. We quantify the significance of the correlation using MC simulations. We find that the subsample of TDEs with IR echoes presents the most significant correlation with IceCube neutrinos, though still below the confidence level. We also discuss the case that includes a systematic error in the events’ error radii. Given that only the statistical error is publicly available for alert events, we add a systematic error of , which is the upper limit of the systematic error estimated by Aartsen et al. (2013c), in quadrature to the statistical error. There are no additional associated events arising from the inclusion of systematic error and the obtained results could be treated as the upper limits of the values. Furthermore, for the potentially associated events reported so far, the arrival dates of neutrinos are within an interval of year after the discovery dates of the TDEs/TDE candidates. Therefore, we tentatively introduce a time delay () term into the weighting factor in our MC simulation to reduce the weight of large time-delayed events. By combining the time delay factor, the best values decrease to 0.20 and 0.0077 () for the all-TDE sample and the subsample with IR echoes, respectively. However, the lowest value remains below the 3 confidence level, even though trial factors have not been taken into account yet.
@@
In total, at present, no significant correlation is found in the population analysis between the ZTF-BTS TDEs and IceCube alert neutrino events. Our work is the first research to quantify the -value of the TDE-neutrino association with actual TDE and neutrino observations. We find that the TDEs with IR flares have a stronger correlation with high-energy neutrinos, consistent with previous findings. Another finding of this work is the discovery of a new TDE–neutrino association: AT2019azh and IC230217A. Although IC230217A occurred a relatively long time after the IR/optical peak of AT2019azh, the TDE exhibits unusually long-lived radio and X-ray emission, making it still plausible that IC230217A originated from AT2019azh. This represents another neutrino alert discovered to be associated with a TDE.
@@
In our sample, only seven TDEs have IR flares and two of them satisfy the spatial-temporal association with high-energy neutrinos. The number of TDEs with IR flares is still limited, which hampers us from obtaining conclusive results. In the future, the persistent optical survey conducted by ZTF-BTS (Perley et al., 2020) is expected to detect more TDEs, in the order of 10 TDEs per year. Moreover, hundreds to thousands of TDEs are expected to be discovered per year by advanced surveys such as the Legacy Survey of Space and Time (LSST) at the Vera Rubin Observatory (Ivezić et al., 2019; Bricman & Gomboc, 2020) and the Wide Field Survey Telescope (WFST, Lin et al. (2022); Wang et al. (2023)). It is also expected that the planned NEO Surveyor (Mainzer et al., 2023), the successor to the NEOWISE mission, will detect more IR flares and help build a larger sample of TDEs with IR echoes, addressing the current issue of a too small sample size. Furthermore, neutrino events with better signalness and directional resolution are also crucial to solve the TDE-neutrino correlation mystery. The next-generation neutrino observatories would also play important roles, such as IceCube-Gen2 (Aartsen et al., 2014c), KM3NeT (Adrian-Martinez et al., 2016b), GRAND (Fang et al., 2018), Baikal-GVD (Avrorin et al., 2018), and RNO-G (Aguilar et al., 2021). By identifying more cases of associations between IR-TDEs and high-energy neutrinos, we will hopefully further uncover the intrinsic connection between the TDE population and high-energy neutrinos, and understand the role that IR echoes play in the neutrino emission process.
On the very day we were about to submit our reply to the second-round review, we noticed that an IceCube Collaboration’s proceeding for ICRC 2025 appeared online (Zegarelli et al., 2025). This proceeding pointed out that the updated version of the alert data has smaller error circles, which places AT2019dsg, AT2019fdr and AT2019aalc outside the error circles, and therefore no longer associated with the neutrino events. This challenges the possibility that these sources are potential neutrino emitters. However, since the IceCat-2 data has not yet been officially released, we will leave the study of their association with TDEs for future work. If a certain level of systematic uncertainty is considered, as discussed in Sec. 4.2, the analysis and conclusions presented in this paper will remain applicable.